# QuiZSF: An efficient data-model interaction framework for zero-shot time-series forecasting

### v1

@@ 5.5 Detailed model analysis
Figure 4: (a) Size and time overview of QuiZTST vs. pre-trained TS benchmarks. Plot each model based on model size and the CPU inference time per batch. (b) Hyperparameter analysis.
Case study. We visualize results on the ETTh2 dataset in Figure 5. The retrieved patterns help smooth out noisy predictions and reduce hallucinations. A complete analysis of sequence selection and LLM behavior is included in Appendix D.3.
Figure 5: Case studies on ETTh2 prediction.
6 Conclusion
We present QuiZSF, a lightweight and efficient framework that integrates TSPMs with RAG for zero-shot time series forecasting. By combining tree-structured base construction, hierarchical retrieval, multi-level representation learning, and modality-aware integration, QuiZSF effectively utilizes both inter- and intra-domain sequences to enhance forecasting accuracy. QuiZSF achieves strong performance in both zero-shot and full-shot settings and offers practical insights for time series modeling in data-scarce scenarios such as domain transfer and extreme conditions. It also holds promise for broader applications in intelligent data systems and temporal databases.
@@ Appendix A Detailed Design of ChronoRAG Base
A.1 Data protocol for ChronoRAG
Large-scale time series datasets are essential for retrieval tasks, but constructing the ChronoRAG Base (CRB) presents challenges such as inconsistent lengths, dimensionality mismatches, missing values, metadata diversity, and storage scalability. To tackle these issues, we design a unified data protocol that ensures consistent preprocessing, metadata unification, and efficient storage. All data is stored using the ARROW format peltenburg2019fletcher , which is optimized for deep learning frameworks and enables efficient retrieval and access.
To address the issue of inconsistent data lengths across datasets, we adopt a sliding window approach. Let $\mathbf{X}\in\mathbb{R}^{N\times D}$ be a multivariate time series with $N$ time steps and $D$ channels. The $j$-th channel is denoted as $\mathbf{x}_{j}=[x_{1j},x_{2j},\cdots,x_{Nj}]$. Following the foundation model settings ekambaram2024tiny , we set window size $w$ and step size $s$ ($w\leq N$), and segment each channel into uniform windows:
$\mathbf{X}_{kj}=[x_{(k-1)s+1,j},\cdots,x_{(k-1)s+w,j}]$ (15)
@@ A.1 Data protocol for ChronoRAG
To address the issue of varying dimensionality, we use a channel-independent strategy, which treats each dimension separately and has been validated by PatchTST nie2022time , TimeLLM jin2023time , and TTM ekambaram2024tiny . For each channel $\mathbf{x}_{i}$, we apply a shared function $f$:
$\mathbf{y}_{i}=f(\mathbf{x}_{i})$ (16)
@@
To address the issue of missing values, which may impair data integrity and affect retrieval, we apply linear interpolation friedman1962interpolation to complete incomplete sequences.
@@
To address the issue of large volume and variety of sequence data, which make efficiently storing and retrieving a significant challenge, we implement a hierarchical tree-like storage structure berchtold1996x ; jensen2004query ; athanassoulis2014bf . This enables efficient storage and indexing for large-scale datasets and seamless integration into deep learning frameworks. Details are provided in Section 6.
Table 4: Structural key-value instance in CRB
@@
Key
Domain
Category
Item_id Start End Freq Target
@@
A.2 Hierarchical Series Tree
Figure 6: Tree-shaped hierarchical series organization with pre-clustering. Prototypes represent cluster centroids.
To support efficient indexing and retrieval in ChronoRAG Base (CRB), we design a hierarchical tree structure with pre-clustering. Traditional FIFO-based linear storage pelkonen2015gorilla ; jensen2017time suffers from inefficiency when scaling to millions of time series. Linear retrieval requires one-by-one comparisons with time complexity $T_{\text{linear}}=O(N)$, which becomes prohibitive at large scale.
To mitigate this, we construct a tree-shaped structure inspired by database indexing techniques berchtold1996x ; jensen2004query ; athanassoulis2014bf . At the top level, the database is partitioned by domain: given a dataset $\mathcal{X}=\{X_{1},\dots,X_{N}\}$, we separate it into $K$ disjoint domain-based groups $\{\mathcal{D}_{k}\}_{k=1}^{K}$, such that:
$\mathcal{X}=\bigcup_{k=1}^{K}\mathcal{D}_{k},\quad\mathcal{D}_{i}\cap\mathcal{D}_{j}=\emptyset\text{ for }i\neq j.$ (17)
Each domain group $\mathcal{D}_{k}$ is then recursively divided using the $k$-means algorithm, where each cluster contains at most $N=256$ time series. This setting follows the clustering granularity used in prior work such as Marigold mortensen2023marigold , where small prototype groups (typically $N\leq 256$) are shown to improve retrieval quality and update flexibility. The number of clusters $M_{k}$ in each domain is thus determined by the size of $\mathcal{D}_{k}$, i.e., $M_{k}=\lceil\frac{|\mathcal{D}_{k}|}{N}\rceil$.
Each cluster $\mathcal{C}_{m}^{(k)}$ is formed by minimizing the standard $k$-means objective:
$\mathcal{L}_{k}=\sum_{X_{i}\in\mathcal{D}_{k}}\|X_{i}-C_{j(i)}^{(k)}\|^{2},$ (18)
where $C_{j(i)}^{(k)}$ is the centroid of the cluster to which $X_{i}$ belongs. The prototype for each cluster is selected as the sequence closest to the centroid:
$X_{\text{proto}}^{(k,m)}=\arg\min_{X_{i}\in\mathcal{C}_{m}^{(k)}}\|X_{i}-C_{m}^{(k)}\|^{2}.$ (19)
While tree-based structures have a theoretical average-case complexity of $O(\log_{b}N)$, this does not always hold in high-dimensional time series due to the curse of dimensionality beyer1999nearest . Hence, instead of relying solely on theoretical claims, we report empirical improvements in retrieval speed and memory in Section 5.5.
@@ A.2 Hierarchical Series Tree
$X_{\text{proto}}^{*}=\arg\min_{X_{\text{proto}}}\|X_{\text{new}}-X_{\text{proto}}\|^{2},$ (20)
@@
$\min_{\{\mathcal{C}_{i}^{(k)}\}}\sum_{X_{j}\in\bigcup\mathcal{C}_{i}^{(k)}}\left\|X_{j}-C_{i(j)}^{(k)}\right\|^{2},$ (21)
@@
The above update strategy allows the index to evolve incrementally over time without full reorganization, and is inspired by dynamic clustering methods in data streams cao2006density ; aghabozorgi2015time .
A.3 Database Versions
The three versions of CRB are illustrated in Figure 7. Each version covers all seven domains, and the smaller versions are subsets of the larger ones. Detailed information about CRB-Large can be found in Table 5.
Figure 7: Three versions of CRB.
Table 5: CRB-Large Detailed Descriptions: Domain indicates the field to which the dataset belongs. Datasets refer to the specific datasets included. Time Series represents the number of time series contained in the dataset after processing. Frequency denotes the sampling interval of time points, where “-” indicates either the absence of timestamps or irregular intervals. Time Points represents the total number of time points in the dataset. Source specifies the original paper or resource from which the dataset is obtained.
Domain Datasets Time Series Frequency Time Points Source
Energy wind_4_seconds_dataset 1 4 Sec 512 Monash godahewa2021monash
australian_electricity_demand_dataset 5 30 Min 2560 Monash godahewa2021monash
london_smart_meters_dataset_without_missing_values 5556 Hourly 2844672 Monash godahewa2021monash
Health SelfRegulationSCP1 3366 0.004 Sec 1723392 UCR Time Series Archive dau2019ucr
MotorImagery 24192 0.001 Sec 12386304 UCR Time Series Archive dau2019ucr
PigCVP 312 - 159744 UCR Time Series Archive dau2019ucr
PigArtPressure 312 - 159744 UCR Time Series Archive dau2019ucr
SelfRegulationSCP2 2660 0.004 Sec 1361920 UCR Time Series Archive dau2019ucr
AtrialFibrillation 60 0.008 Sec 30720 UCR Time Series Archive dau2019ucr
IEEEPPG 15480 0.008 Sec 7925760 TSER archive tan2021time
BIDMC32HR 12278 - 6286336 TSER archive tan2021time
TDBrain 28644 0.002 Sec 14665728 TDBrain van2022two
IoT baian 918 0.02 Sec 470016 UTSD  liutimer
Nature StarLightCurves 9236 - 4728832 UCR Time Series Archive dau2019ucr
Phoneme 2110 - 1080320 UCR Time Series Archive dau2019ucr
EigenWorms 1554 - 795648 UCR Time Series Archive dau2019ucr
Worms 258 0.033 Sec 132096 UCR Time Series Archive dau2019ucr
us_births_dataset 1 Daily 512 Monash godahewa2021monash
kdd_cup_2018_dataset_without_missing_values 270 Hourly 138240 Monash godahewa2021monash
temperature_rain_dataset_without_missing_values 32072 Daily 16420864 Monash godahewa2021monash
Sunspot_dataset_without_missing_values 1 Daily 512 Monash godahewa2021monash
saugeenday_dataset 1 Daily 512 Monash godahewa2021monash
Web kaggle_web_traffic_dataset_without_missing_values 141444 Daily 72419328 Monash godahewa2021monash
Transport pedestrian_counts_dataset 66 Hourly 33792 Monash godahewa2021monash
Environment AustraliaRainfall 3 Hourly 1536 TSER archive tan2021time
BenzeneConcentration 8 Hourly 4096 TSER archive tan2021time
BeijingPM25Quality 9 Hourly 4608 TSER archive tan2021time
@@ Appendix B Language Coherer for LLM based TSPMs
Unlike numerical models, language models operate exclusively on text. Direct integration of numeric features is thus infeasible due to modality mismatch. To bridge this gap, we transform the MSIL-derived representations (i.e., $\mathbf{P}_{\text{int}}$, $\mathbf{P}_{\text{avg}}$, and $\mathbf{T}^{\text{norm}}$) into structured textual summaries. These summaries are combined with an instruction prompt to guide the language models in generating forecasting outputs.
@@
Figure 8: Prompt construction framework for LLM-based forecasting.
@@ Appendix C Training Dataset Details
Figure 9: Detailed information of the training set.
@@ D.1 Complexity Analysis
D.2 Hyperparameter Study and Analysis
@@
The CRB_Small scale performs worst (Figure 10 (a)). As the scale decreases, the external knowledge it provides declines, leading to poor performance. This partly verifies the scaling law kaplan2020scaling in time series. In the hyperparameter experiment for the retrieved number K, TopK =8=8 yields the best results. Retrieving more sequences may introduce more noise, while fewer sequences carry less information (Figure 10 (b)). In the hyperparameter experiment for the Local prototype $\rho$, two tasks show different trends. A too-high proportion of the same domain limits data feature diversity, over-emphasizing single-dimension features. The model performs well on samples fitting this feature (Figure 10 (c)), but poorly on non-matching samples due to the lack of auxiliary correction from other dimensions, resulting in a polarized outcome (Figure 10 (d)). To balance performance and efficiency, we select CRB_Medium, set TopKK to 8, and set the Local prototype $\rho$ at $60\%$.
Figure 10: Hyperparameter analysis.
D.3 Case Study
@@
D.4 Experiments compute resources
@@ Appendix E Limitations and Future Work
While our proposed framework, QuiZSF, demonstrates significant advancements in zero-shot forecasting for time series data, there are several limitations that warrant further investigation. First, the effectiveness of QuiZSF is tied to the quality and relevance of pre-trained models, which can be a limiting factor in domains with limited or highly specialized data. Future work could explore the development of more robust and domain-specific pre-trained models to enhance the framework’s capabilities. Second, reducing the complexity of the framework is crucial for large-scale retrieval processes. Introducing sparse connections into the retrieval database and exploring such connections to facilitate relevance retrieval could significantly improve efficiency. Lastly, developing multi-granularity sequential learning methods to allow pattern transfer between different sub-sequences could address cross-granularity transfer challenges. These advancements would further enhance the scalability and applicability of QuiZSF in diverse forecasting scenarios.

### v2

@@ 5.5. Detailed model analysis
5.6. Case study
6. Conclusion
We presented QuiZSF, a retrieval-augmented framework for zero-shot time series forecasting that unifies search and forecasting in dynamic Web environments. QuiZSF combines a tree-structured temporal database, hybrid retrieval strategies, multi-level interaction learning, and modality-aware adaptation, enabling models to retrieve auxiliary sequences and incorporate them for more accurate prediction. Experiments on public benchmarks show that QuiZSF achieves state-of-the-art performance with both Non-LLM and LLM-based TSPMs, while maintaining efficiency in memory and inference. Beyond forecasting, this work illustrates how retrieval-augmented AI can extend beyond text to time series, offering new perspectives for the WWW community and opening opportunities for adaptive and intelligent Web-scale systems.
7. acknowledgements
This paper is partially supported by the National Natural Science Foundation of China (No.62502488, No.12227901), Natural Science Foundation of Jiangsu Province (BK20240460), the grant from State Key Laboratory of Resources and Environmental Information System.
@@ Appendix A Detailed Design of ChronoRAG Base
A.1. Data protocol for ChronoRAG
Large-scale time series datasets are essential for retrieval tasks, but constructing the ChronoRAG Base (CRB) presents challenges such as inconsistent lengths, dimensionality mismatches, missing values, metadata diversity, and storage scalability. To tackle these issues, we design a unified data protocol that ensures consistent preprocessing, metadata unification, and efficient storage. All data is stored using the ARROW format (Peltenburg et al., 2019), which is optimized for deep learning frameworks and enables efficient retrieval and access.
To address the issue of data length inconsistencies across datasets, we adopt a sliding window approach. Let $\mathbf{X}\in\mathbb{R}^{N\times D}$ be a multivariate time series with $N$ time steps and $D$ channels. The $j$-th channel is denoted as $\mathbf{x}_{j}=[x_{1j},x_{2j},\cdots,x_{Nj}]$. Following the foundation model settings (Ekambaram et al., 2024), we set window size $w$ and step size $s$ ($w\leq N$), and segment each channel into uniform windows:
(13) $\mathbf{X}_{kj}=[x_{(k-1)s+1,j},\cdots,x_{(k-1)s+w,j}]$
@@ A.1. Data protocol for ChronoRAG
To address the issue of varying dimensionality, we use a channel-independent strategy, which treats each dimension separately and has been validated by PatchTST (Nie et al., 2022), TimeLLM (Jin et al., 2023), and TTM (Ekambaram et al., 2024). For each channel $\mathbf{x}_{i}$, we apply a shared function $f$:
(14) $\mathbf{y}_{i}=f(\mathbf{x}_{i})$
@@
To address the issue of missing values, which may impair data integrity and affect retrieval, we apply linear interpolation (Friedman, 1962) to complete incomplete sequences.
@@
To address the issue of large volume and variety of sequence data, which make efficiently storing and retrieving a significant challenge, we implement a hierarchical tree-like storage structure (Berchtold et al., 1996; Jensen et al., 2004; Athanassoulis and Ailamaki, 2014). This enables efficient storage and indexing for large-scale datasets and seamless integration into deep learning frameworks. Details are provided in Section 7.
Table 5. Structural key-value instance in CRB
@@
Key Domain Category Item_id Start End Freq Target
@@
A.2. Hierarchical Series Tree
To support efficient indexing and retrieval in ChronoRAG Base (CRB), we design a hierarchical tree structure with pre-clustering. Traditional FIFO-based linear storage (Pelkonen et al., 2015; Jensen et al., 2017) suffers from inefficiency when scaling to millions of time series. Linear retrieval requires one-by-one comparisons with time complexity $T_{\text{linear}}=O(N)$, which becomes prohibitive at large scale.
To mitigate this, we construct a tree-shaped structure inspired by database indexing techniques (Berchtold et al., 1996; Jensen et al., 2004; Athanassoulis and Ailamaki, 2014). At the top level of the structure, the database is partitioned by domain: given a dataset $\mathcal{X}=\{X_{1},\dots,X_{N}\}$, we separate it into $K$ disjoint domain-based groups $\{\mathcal{D}_{k}\}_{k=1}^{K}$, such that:
(15) $\mathcal{X}=\bigcup_{k=1}^{K}\mathcal{D}_{k},\quad\mathcal{D}_{i}\cap\mathcal{D}_{j}=\emptyset\text{ for }i\neq j.$
Each domain group $\mathcal{D}_{k}$ is then recursively divided using the $k$-means algorithm, where each cluster contains at most $N=256$ time series. This setting follows the clustering granularity used in prior work such as Marigold (Mortensen et al., 2023), where small prototype groups (typically $N\leq 256$) are shown to improve retrieval quality and update flexibility. The number of clusters $M_{k}$ in each domain is thus determined by the size of $\mathcal{D}_{k}$, i.e., $M_{k}=\lceil\frac{|\mathcal{D}_{k}|}{N}\rceil$.
Each cluster $\mathcal{C}_{m}^{(k)}$ is formed by minimizing the standard $k$-means objective function:
(16) $\mathcal{L}_{k}=\sum_{X_{i}\in\mathcal{D}_{k}}\|X_{i}-C_{j(i)}^{(k)}\|^{2},$
where $C_{j(i)}^{(k)}$ is the centroid of the cluster to which $X_{i}$ belongs. For each cluster, the prototype is selected as the sequence closest to its centroid:
(17) $X_{\text{proto}}^{(k,m)}=\arg\min_{X_{i}\in\mathcal{C}_{m}^{(k)}}\|X_{i}-C_{m}^{(k)}\|^{2}.$
While tree-based structures have a theoretical average-case complexity of $O(\log_{b}N)$, this does not always hold in high-dimensional time series due to the curse of dimensionality (Beyer et al., 1999). Hence, instead of relying solely on theoretical claims, we report empirical improvements in retrieval speed and memory in Section 5.5.
@@ A.2. Hierarchical Series Tree
(18) $X_{\text{proto}}^{*}=\arg\min_{X_{\text{proto}}}\|X_{\text{new}}-X_{\text{proto}}\|^{2},$
@@
(19) $\min_{\{\mathcal{C}_{i}^{(k)}\}}\sum_{X_{j}\in\bigcup\mathcal{C}_{i}^{(k)}}\left\|X_{j}-C_{i(j)}^{(k)}\right\|^{2},$
@@
The above update strategy allows the index to evolve incrementally over time without full reorganization, and is inspired by dynamic clustering methods in data streams (Cao et al., 2006; Aghabozorgi et al., 2015).
Figure 7. Tree-shaped hierarchical series organization with pre-clustering. Prototypes represent cluster centroids.
@@ Appendix B Language Coherer for LLM based TSPMs
Unlike numerical models, language models operate exclusively on text. Direct integration of numeric features is thus infeasible due to modality mismatch. To bridge this gap, we transform the MSIL-derived representations (i.e., $\mathbf{P}_{\text{int}}$, $\mathbf{P}_{\text{avg}}$, and $\mathbf{T}^{\text{norm}}$) into structured textual summaries (see Figure 9). These summaries are combined with an instruction prompt to guide the language models in generating forecasting outputs.
@@ Appendix C Training Dataset Details
Figure 8. Detailed information of the training set.
@@ D.1. Complexity Analysis
Figure 9. Prompt construction framework for LLM-based forecasting.
D.2. Hyperparameter Study and Analysis
@@
The CRB_Small scale performs worst (Figure 10 (a)). As the scale decreases, the external knowledge it provides declines, leading to poor performance. This partly verifies the scaling law (Kaplan et al., 2020) in time series. In the hyperparameter experiment for the retrieved number K, TopK =8=8 yields the best results. Retrieving more sequences may introduce more noise, while fewer sequences carry less information (Figure 10 (b)). In the hyperparameter experiment for the Local prototype $\rho$, two tasks show different trends. A too-high proportion of the same domain limits data feature diversity, over-emphasizing single-dimension features. The model performs well on samples fitting this feature (Figure 10 (c)), but poorly on non-matching samples due to the lack of auxiliary correction from other dimensions, resulting in a polarized outcome (Figure 10 (d)). To balance performance and efficiency, we select CRB_Medium, set TopKK to 8, and set the Local prototype $\rho$ at $60\%$.
Figure 10. Hyperparameter analysis.
D.3. Case Study
@@
D.4. Experiments compute resources
@@ Appendix E Limitations and Future Work
QuiZSF, while advancing zero-shot time series forecasting, faces limitations: its effectiveness relies on pre-trained model quality, which is limited in data-scarce domains; large-scale retrieval efficiency remains a challenge, potentially addressable with sparse indexing; and multi-granularity sequence learning for cross-pattern transfer is needed to enhance generalization across heterogeneous streams, ultimately strengthening QuiZSF and contributing to scalable retrieval-augmented AI systems for diverse Web environments.